# Offline-online approximation of multiscale eigenvalue problems with random defects

### v1

@@ 3 Preliminaries for the error analysis
$E_{l}(\lambda^{(l)}):=\text{span}\{u^{(1)},\cdots,u^{(l)}\}.$
@@
Find $Tf\in\mathcal{H}^{1}_{\#}$ such that
$a(Tf,v)=m(f,v),\;\;\forall v\in\mathcal{H}^{1}_{\#}.$ (19)
Find $T^{1}_{H}f\in V_{H}\subset\mathcal{H}^{1}_{\#}$ such that
$a_{1}(T^{1}_{H}f,v)=m_{1}(f,v),\;\;\forall v\in V_{H}.$ (20)
@@
$a_{2}(T^{2}_{H}f,v)=m_{2}(f,v),\;\;\forall v\in V_{H}.$ (21)
For $f\in V_{H}$, the function $T^{2}_{H}f$ is defined and is taken as a map on $V_{H}$. While $T$ is viewed as an operator on #$\mathcal{H}^{1}_{\#}$, the operator $T^{1}_{H}$ can be considered either to be on #$\mathcal{H}^{1}_{\#}$ or on $V_{H}$. For the perturbation analysis, we need to relate the error $T^{1}_{H}-T^{2}_{H}$ to the errors in $m_{1}-m_{2}$ and $a_{1}-a_{2}$ in the following.
@@
$\begin{split}\hskip 56.9055pt\norm{T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)}}^{2}_{\mathcal{A}}&\lesssim\absolutevalue{(m_{1}-m_{2})(f^{(l+j)},T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)})}\\ &\hskip 42.67912pt+\absolutevalue{(a_{2}-a_{1})(T^{2}f^{(l+j)},T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)})}.\end{split}$
@@
$\norm{T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)}}^{2}_{\mathcal{A}}$ $\lesssim a_{1}(T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)},T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)})$
$=a_{1}(T^{1}_{H}f^{(l+j)},T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)})-a_{2}(T^{2}_{H}f^{(l+j)},T^{2}_{H}f^{(l+j)}-T^{1}_{H}f^{(l+j)})$
$\hskip 34.14322pt+a_{2}(T^{2}_{H}f^{(l+j)},T^{2}_{H}f^{(l+j)}-T^{1}_{H}f^{(l+j)})-a_{1}(T^{2}_{H}f^{(l+j)},T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)})$
|$\leq\absolutevalue{(m_{1}-m_{2})(f^{(l+j)},T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)})}$ |
+|$\hskip 34.99677pt+\absolutevalue{(a_{2}-a_{1})(T^{2}f^{(l+j)},T^{1}_{H}f^{(l+j)}-T^{2}_{H}f^{(l+j)})}.\qed$
@@ 4 Error analysis
In this section, we derive error estimates for eigenfunctions and eigenvalues for the MLOD and the OLOD method. Recall that in the MLOD formulation, we have considered the trial space to be the multiscale space and the test space to be the classical finite element space on the coarse mesh. Consequently, the stiffness and mass matrix in the MLOD are based on the coarse FE space and different from the ones used in the Galerkin LOD method in [MP15]. We therefore start with several preliminary results particularly estimating the errors from these changes. Note that our final goal are estimates for the OLOD method, for which – similar to [BO90] — the MLOD method is introduced and analyzed as an intermediate step.
@@ 4.1 Preliminary estimates for the MLOD
For any $v_{1},v_{2}\in V=\mathcal{H}^{1}_{\#}$, it holds that
|$\begin{split}\frac{\absolutevalue{({m}_{\mathfrak{k}}-\widetilde{m}_{\mathfrak{k}})(v_{1},v_{2})}}{\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}}&\lesssim H^{2}.\end{split}$
@@
|$\begin{split}\absolutevalue{({m}_{\mathfrak{k}}-\widetilde{m}_{\mathfrak{k}})(v_{1},v_{2})}&\lesssim H^{2}\Big{(}\norm{v_{1}}_{\mathcal{A}}\norm{\mathscr{C}_{\mathfrak{k}}v_{2}}_{\mathcal{A}}+\norm{v_{2}}_{\mathcal{A}}\norm{\mathscr{C}_{\mathfrak{k}}v_{1}}_{\mathcal{A}}+\norm{\mathscr{C}_{\mathfrak{k}}v_{1}}_{\mathcal{A}}\norm{\mathscr{C}_{\mathfrak{k}}v_{2}}_{\mathcal{A}}\Big{)}\\ &\leq H^{2}\Bigg{[}\norm{v_{1}}_{\mathcal{A}}\Big{(}\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}+\norm{\mathscr{C}v_{2}}_{\mathcal{A}}\Big{)}+\norm{v_{2}}_{\mathcal{A}}\Big{(}\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1}}_{\mathcal{A}}+\norm{\mathscr{C}v_{1}}_{\mathcal{A}}\Big{)}\\ &\hskip 136.5733pt+\big{(}\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1}}_{\mathcal{A}}+\norm{\mathscr{C}v_{1}}_{\mathcal{A}}\big{)}\big{(}\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}+\norm{\mathscr{C}v_{2}}_{\mathcal{A}}\big{)}\Bigg{]}\\ &\lesssim H^{2}\Big{[}3+4g(\mathfrak{k})+g^{2}(\mathfrak{k})\Big{]}\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}.\end{split}$
@@
Similar in the previous lemma, we use the definitions of $\overline{a}(\cdot,\cdot)$ and $\widetilde{a}(\cdot,\cdot)$ to obtain
$\begin{split}({a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}})(v_{1},v_{2})&=a((1-\mathscr{C}_{\mathfrak{k}})v_{1},(1-\mathscr{C}_{\mathfrak{k}})v_{2})-a((1-\mathscr{C}_{\mathfrak{k}})v_{1},v_{2})\\ &=a((1-\mathscr{C}_{\mathfrak{k}})v_{1},-\mathscr{C}_{\mathfrak{k}}v_{2})\\ &=a((1-\mathscr{C}_{\mathfrak{k}})v_{1},(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2})-a((1-\mathscr{C}_{\mathfrak{k}})v_{1},\mathscr{C}v_{2})\\ &=a(v_{1},(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2})-a(\mathscr{C}_{\mathfrak{k}}v_{1},(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}).\end{split}$
This implies together with the stability of $\mathscr{C}$ in the energy norm that
$\begin{split}({a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}})(v_{1},v_{2})&\lesssim\norm{v_{1}}_{\mathcal{A}}\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}+\norm{\mathscr{C}_{\mathfrak{k}}v_{1}}_{\mathcal{A}}\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}\\ &\lesssim\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}\Big{(}\norm{v_{1}}_{\mathcal{A}}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1}}_{\mathcal{A}}+\norm{\mathscr{C}v_{1}}_{\mathcal{A}}\Big{)}\\ &\lesssim g(\mathfrak{k})\Big{(}2+g(\mathfrak{k})\Big{)}\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}.\end{split}$
@@
If taken $\mathfrak{k}$ large enough such that $g(\mathfrak{k})=H^{2}$, Lemmas 4.1 and 4.2 tells us that the error between mass matrices due to the selection of the FEM mass matrix is the same as the error due to the Petrov-Galerkin localization in the stiffness matrix.
@@
Suppose $(\lambda,u)$ is an eigenpair of (2). Further let, $\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}_{L^{2}(\Omega)}=1$ with $v^{\mathrm{ms}}_{H,\mathfrak{k}},=v_{H,\mathfrak{k}}-\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}}$ and $\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\neq 0$ for $v_{H,\mathfrak{k}}\in V_{H}\subset\mathcal{H}^{1}_{\#}$ and $v^{\mathrm{ms}}_{H,\mathfrak{k}}\in V^{\mathrm{ms}}_{H,\mathfrak{k}}\subset\mathcal{H}^{1}_{\#}$ Consider the MLOD formulation of (2) given in (9). Then,
$\begin{split}\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}-\lambda\leq&\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 56.9055pt+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\Big{(}{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\Big{)}.\end{split}$
@@
$\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}=\norm{u}^{2}_{\mathcal{A}}-2a(u,v^{\mathrm{ms}}_{H,\mathfrak{k}})+\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u}_{L^{2}(\Omega)}+2\lambda m(u,v^{\mathrm{ms}}_{H,\mathfrak{k}})-\lambda\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}.$ (22)
@@
Recall that $\widetilde{a}_{\mathfrak{k}}(\cdot,\cdot)=a((\cdot)^{\mathrm{ms}}_{\mathfrak{k}},(\cdot))$ and $\widetilde{m}_{\mathfrak{k}}(\cdot,\cdot)=m((\cdot),(\cdot))=\norm{(\cdot)}^{2}_{L^{2}(\Omega)}$. Then we deduce that
$\begin{split}a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v^{\mathrm{ms}}_{H,\mathfrak{k}})&=a(v^{\mathrm{ms}}_{H,\mathfrak{k}},(1-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}})\\ &=a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-a((1-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}},\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}})\\ &=a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-a(v_{H,\mathfrak{k}}-\mathscr{C}v_{H,\mathfrak{k}},\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}})-a((\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}},\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}}).\end{split}$
Due to the $a$-orthogonality, and the stability of $\mathscr{C}_{\mathfrak{k}}$ in the energy norm we derive
$\begin{split}a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}})&=a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v^{\mathrm{ms}}_{H,\mathfrak{k}})+a((\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}},\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}})\\ &\leq a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v^{\mathrm{ms}}_{H,\mathfrak{k}})+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$ (24)

### v2

@@ 3 Preliminaries for the error analysis
${E^{(l)}}(\lambda^{(l)}):=\text{span}\left\{u^{(1)},\cdots,u^{(l)}\right\}.$
@@
Find $Tf\in\mathcal{H}^{1}_{\#}(\Omega)$ such that
$a(Tf,v)=m(f,v),\;\;\forall v\in\mathcal{H}^{1}_{\#}(\Omega).$
Find $T^{1}_{H}f\in V_{H}\subset\mathcal{H}^{1}_{\#}(\Omega)$ such that
$a_{1}\left(T^{1}_{H}f,v\right)=m_{1}\left(f,v\right),\;\;\forall v\in V_{H}.$ (17)
@@
$a_{2}\left(T^{2}_{H}f,v\right)=m_{2}\left(f,v\right),\;\;\forall v\in V_{H}.$ (18)
In all cases above $f\in\mathcal{H}^{1}_{\#}(\Omega)$. For $f\in V_{H}$, the function $T^{2}_{H}f$ is defined and is taken as a map on $V_{H}$. While $T$ is viewed as an operator on #$\mathcal{H}^{1}_{\#}(\Omega)$, the operator $T^{1}_{H}$ can be considered either to be on #$\mathcal{H}^{1}_{\#}(\Omega)$ or on $V_{H}$. For the perturbation analysis, we need to relate the error $T^{1}_{H}-T^{2}_{H}$ to the errors in $m_{1}-m_{2}$ and $a_{1}-a_{2}$ in the following.
@@
$\begin{split}\hskip 56.9055pt\norm{T^{1}_{H}f-T^{2}_{H}f}^{2}_{\mathcal{A}}&\lesssim\absolutevalue{\left(m_{1}-m_{2}\right)\left(f,T^{1}_{H}f-T^{2}_{H}f\right)}\\ &\hskip 42.67912pt+\absolutevalue{\left(a_{2}-a_{1}\right)\left(T^{2}_{H}f,T^{1}_{H}f-T^{2}_{H}f\right)}.\end{split}$
@@
$\norm{T^{1}_{H}f-T^{2}_{H}f}^{2}_{\mathcal{A}}$ $\lesssim a_{1}\left(T^{1}_{H}f-T^{2}_{H}f,T^{1}_{H}f-T^{2}_{H}f\right)$
$=a_{1}\left(T^{1}_{H}f,T^{1}_{H}f-T^{2}_{H}f\right)-a_{2}\left(T^{2}_{H}f,T^{1}_{H}f-T^{2}_{H}f\right)$
$\hskip 56.9055pt+a_{2}\left(T^{2}_{H}f,T^{1}_{H}f-T^{2}_{H}f\right)-a_{1}\left(T^{2}_{H}f,T^{1}_{H}f-T^{2}_{H}f\right)$
|$\leq\absolutevalue{(m_{1}-m_{2})(f,T^{1}_{H}f-T^{2}_{H}f)}+\absolutevalue{(a_{2}-a_{1})(T^{2}_{H}f,T^{1}_{H}f-T^{2}_{H}f)}.\qed$
@@ 4 Error analysis
In this section, we derive error estimates for eigenfunctions and eigenvalues for the MLOD and the OLOD methods. Recall that in the MLOD formulation, we have considered the trial space to be the multiscale space and the test space to be the classical finite element space on the coarse mesh. Consequently, the stiffness and mass matrix in the MLOD are based on the coarse FE space and different from the ones used in the Galerkin LOD method in [MP15]. We therefore start with several preliminary results particularly estimating the errors from these changes. Note that our final goal is to obtain estimates for the OLOD method, for which – similar to [BO90] — the MLOD method is introduced and analyzed as an intermediate step.
@@ 4.1 Preliminary estimates for the MLOD
For any $v_{1},v_{2}\in V=\mathcal{H}^{1}_{\#}(\Omega)$, it holds that
|$\begin{split}\frac{\absolutevalue{\left({m}_{\mathfrak{k}}-\widetilde{m}_{\mathfrak{k}}\right)(v_{1},v_{2})}}{\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}}&\lesssim H^{2}.\end{split}$
@@
|$\begin{split}\absolutevalue{\left({m}_{\mathfrak{k}}-\widetilde{m}_{\mathfrak{k}}\right)(v_{1},v_{2})}&\lesssim H^{2}\left(\norm{v_{1}}_{\mathcal{A}}\norm{\mathscr{C}_{\mathfrak{k}}v_{2}}_{\mathcal{A}}+\norm{v_{2}}_{\mathcal{A}}\norm{\mathscr{C}_{\mathfrak{k}}v_{1}}_{\mathcal{A}}+\norm{\mathscr{C}_{\mathfrak{k}}v_{1}}_{\mathcal{A}}\norm{\mathscr{C}_{\mathfrak{k}}v_{2}}_{\mathcal{A}}\right)\\ &\leq H^{2}\Big{(}\norm{v_{1}}_{\mathcal{A}}\left(\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}+\norm{\mathscr{C}v_{2}}_{\mathcal{A}}\right)+\norm{v_{2}}_{\mathcal{A}}\left(\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1}}_{\mathcal{A}}+\norm{\mathscr{C}v_{1}}_{\mathcal{A}}\right)\\ &\hskip 136.5733pt+\left(\norm{\left(\mathscr{C}-\mathscr{C}_{\mathfrak{k}}\right)v_{1}}_{\mathcal{A}}+\norm{\mathscr{C}v_{1}}_{\mathcal{A}}\right)\left(\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}+\norm{\mathscr{C}v_{2}}_{\mathcal{A}}\right)\Big{)}\\ &\lesssim H^{2}\left(3+4g(\mathfrak{k})+g^{2}(\mathfrak{k})\right)\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}.\end{split}$
@@
We use the definitions of $a_{\mathfrak{k}}\left(\cdot,\cdot\right)$ and $\widetilde{a}_{\mathfrak{k}}(\cdot,\cdot)$ to obtain
$\begin{split}\left({a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}}\right)\left(v_{1},v_{2}\right)&=a\left((1-\mathscr{C}_{\mathfrak{k}})v_{1},(1-\mathscr{C}_{\mathfrak{k}})v_{2}\right)-a\left((1-\mathscr{C}_{\mathfrak{k}})v_{1},v_{2}\right)\\ &=a\left((1-\mathscr{C}_{\mathfrak{k}})v_{1},-\mathscr{C}_{\mathfrak{k}}v_{2}\right)\\ &=a\left((1-\mathscr{C})v_{1}+(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1},-\mathscr{C}_{\mathfrak{k}}v_{2}\right)\\ &=a((\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1},-\mathscr{C}_{\mathfrak{k}}v_{2})\\ &=a((\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1},(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2})+a((\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1},-\mathscr{C}v_{2})\end{split}$
The second last equality follows by the $a$-orthogonality property of the correction operator $\mathscr{C}$. Then by Theorem 2.2 and the stability of $\mathscr{C}$ in the energy norm yields that
$\begin{split}({a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}})(v_{1},v_{2})&\lesssim\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1}}_{\mathcal{A}}\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}\\ &\lesssim\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{1}}_{\mathcal{A}}\left(\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{2}}_{\mathcal{A}}+\norm{v_{2}}_{\mathcal{A}}\right)\\ &\lesssim g(\mathfrak{k})\left(g(\mathfrak{k})+1\right)\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}.\end{split}$
@@
If taken $\mathfrak{k}$ large enough such that $g(\mathfrak{k})\approx H^{2}$, Lemmas 4.1 and 4.2 tells us that the error between mass matrices due to the selection of the FEM mass matrix is the same as the error due to the Petrov-Galerkin localization in the stiffness matrix.
@@
Suppose $(\lambda,u)$ is an eigenpair of (3). Further let, $\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}_{L^{2}(\Omega)}=1$ with $v^{\mathrm{ms}}_{H,\mathfrak{k}},=v_{H,\mathfrak{k}}-\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}}$ and $\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\neq 0$ for $v_{H,\mathfrak{k}}\in V_{H}\subset\mathcal{H}^{1}_{\#}(\Omega)$ and $v^{\mathrm{ms}}_{H,\mathfrak{k}}\in V^{\mathrm{ms}}_{H,\mathfrak{k}}\subset\mathcal{H}^{1}_{\#}(\Omega)$ Consider the MLOD formulation of (3) given in (8). Then,
$\begin{split}\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}-\lambda\leq&\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 56.9055pt+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\left({m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\right).\end{split}$
@@
$\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}=\norm{u}^{2}_{\mathcal{A}}-2a(u,v^{\mathrm{ms}}_{H,\mathfrak{k}})+\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u}_{L^{2}(\Omega)}^{2}+2\lambda m(u,v^{\mathrm{ms}}_{H,\mathfrak{k}})-\lambda\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}.$ (19)
@@
Then we deduce that
$\begin{split}a\left(v^{\mathrm{ms}}_{H,\mathfrak{k}},v^{\mathrm{ms}}_{H,\mathfrak{k}}\right)&=a\left(v^{\mathrm{ms}}_{H,\mathfrak{k}},(1-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}}\right)\\ &=a\left(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}}\right)-a\left((1-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}},\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}}\right)\\ &=a\left(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}}\right)-a\left(v_{H,\mathfrak{k}}-\mathscr{C}v_{H,\mathfrak{k}},\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}}\right)-a\left((\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}},\mathscr{C}_{\mathfrak{k}}v_{H,\mathfrak{k}}\right).\end{split}$
Due to the $a$-orthogonality of $\mathscr{C}$, the middle term vanishes. Then the stability of $\mathscr{C}_{\mathfrak{k}}$ in the energy norm for $v_{H,\mathfrak{k}}\in V_{H}$ yields
$\begin{split}a\left(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}}\right)&\leq a\left(v^{\mathrm{ms}}_{H,\mathfrak{k}},v^{\mathrm{ms}}_{H,\mathfrak{k}}\right)+\norm{\left(\mathscr{C}-\mathscr{C}_{\mathfrak{k}}\right)v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$ (21)